I am stuck on a problem, my teacher only wants me to use the Elimination method
6x-y=20
3x-y=7
I am getting
x=3 but when I substitute for x in the second equation I get
6(3) + y = 20
18 + y = 20
but I can't divide each side by 18 evenly, how do I solve for y?

Answer by ewatrrr(24785)


Hi
solving by Elimination method
6x-y = 20
3x-y = 7
6x-y = 20
-3x+y = -7 | Multiplying the second Eq by -1 to eliminate y
3x = 13
x = 13/3 and y = 6 ( y = 3x -7 from the 2nd origial Eq )
Solution is (13/3, 6)
CHECKING our Answer***
6x-y = 20
6*(13/3)- 6 = 26 - 6 = 20